Assume 2 kg of hydrogen (Item B 28) are initially at 40 atm, 100 K. (a) If the hydrogen changes pressure to 1 atm without a change of temperature, find the changes of enthalpy and entropy. (b) If the hydrogen changes temperature to 35 K with no change of entropy, find the change of pressure and enthalpy.

1
Expert's answer
2021-11-15T10:33:25-0500

a)

"H=\nu c_pT=271 \frac JK,"

"S=\frac{\nu p}{c_pT}=56~\frac JK,"

b)

"p=\frac{\nu (T_2-T_1)}{c_p}=38~Pa,"

"H=\nu pT=71~\frac JK."